hnlxcatherinee:

If the quantity 4 times x times y cubed plus 8 times x squared times y to the fifth power all over 2 times x times y squared is completely simplified to 2xayb + 4xcyd, where a, b, c, and d represent integer exponents, what is the value of a?

Florisalreadytaken:

\[\frac{ (4x \times y^3 + 8 \times x^2 \times y^5) }{ 2x \times y^2 }\]

Florisalreadytaken:

we have got to simplify that into \[2x^ay^b+4x^cy^d\]

Florisalreadytaken:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Florisalreadytaken \[\frac{ (4x \times y^3 + 8 \times x^2 \times y^5) }{ 2x \times y^2 }\] \(\color{#0cbb34}{\text{End of Quote}}\) we can write this equation as \[\frac{ 4xy^3}{2xy^2} + \frac{ 8x^2y^5 }{ 2xy^2 }\]

Florisalreadytaken:

we get rid of the 2s at the bottom, so we have \[\frac{2xy^3 }{ xy^2 }+ \frac{ 4x^2y^5 }{xy^2 }\]

Florisalreadytaken:

thus \[2x^{1-1} y ^{3-2}+4x ^{2-1}y ^{5-2}\]

Florisalreadytaken:

\[2x^0y^1+4x^1y^3\] so we got it in \[2x^ay^b+4x^cy^d\] equation form

Florisalreadytaken:

by comparing them we can obv see that a=0 -- that will be your answer :D